OpenStudy (anonymous):

Find the next term of the following sequence. 25, 10, 4, ... a) 0 b) -5 c) 8/5

OpenStudy (anonymous):

divide by five multiply by 2

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

25/5= 5*2=10 10/5=2*2=4 4/5=4/5*2=8/5